# Deciphering the Factors Influencing the Efficacy of Chain-of-Thought: Probability, Memorization, and Noisy Reasoning

### v1

@@ 4.1 A simple probabilistic approach to modeling the reasoning process
  • $input\_logprob$: log probability of the encoded input text as measured by GPT-2 Radford et al. (2019). The inputs tend to have a very low probability because they are enciphered.
  • $output\_logprob$: log probability of the ground-truth output text as measured by GPT-2.
  • $shift\_freq$: we used the frequency of occurrence of all shift levels that McCoy et al. (2023) provided based on analysis of the C4 corpus Raffel et al. (2020). The assumption is that the distribution of shifts in C4 is similar to the distribution in the training data for GPT-4.
  • $\min(shift\_level,26-shift\_level)$: this value is the minimum number of steps that must be performed to decode each letter, under the assumption that decoding can be done by moving $shift\_level$ steps backward or $(26-shift\_level)$ steps forward; as discussed above, GPT-4 indeed shows evidence of using both of these decoding directions.
@@
$correct\sim\min(shift\_level,26-shift\_level)$
$\qquad\qquad+input\_logprob+output\_logprob$
$\qquad\qquad+shift\_freq$
@@
The following features had a statistically significant effect on model performance: $output\_logprob$, $shift\_frequency$, and $\min(shift\_level,26-shift\_level)$ ($p<10^{-15}$ in all cases). These results therefore quantitatively support the conclusion that GPT-4 incorporates processes based on probability, memorization, and noisy reasoning.
@@
Figure 4 shows the predictions of the logistic regression compared to GPT-4’s actual performance. The logistic regression correctly predicts the main trends in GPT-4’s behavior (as expected, it does not match the curve exactly due to the simplicity of the logistic regression model). The other curves are ablations that use logistic regressions with only one feature among $output\_logprob$, $shift\_freq$, and $shift\_level$. In the next few subsections, we conduct some additional experiments that investigate each hypothesized reasoning type in more detail.
@@ 4.2 Analyzing the effect of probability
If an LLM is influenced by probability, we would expect to occasionally observe unfaithfulness between the chain of reasoning steps produced by the LLM and the LLM’s final answer. Specifically, if the individual reasoning steps would point to a final output that is low probability, a probabilistic reasoner might instead produce a different final answer that is higher probability. For example, in our CoT experiments, each step produces one letter, and these letters must be concatenated to form the final answer. If the individual step outputs are S, T, A, Z, the final answer should be STAZ, but a model might instead “self-correct” by producing the higher probability word STAY.
@@
To investigate unfaithfulness, we compare the faithful accuracy that would be obtained by concatenating the model’s step outputs to the actual overall accuracy. We indeed observe that overall accuracy is generally lower than faithful accuracy, illustrating that unfaithfulness occurs. Further, the drop in accuracy is more pronounced in the low-probability setting than the high-probability setting, which is consistent with the intuition that the lower the probability of a concatenated answer is, the more likely it will be that a probability-reliant model will be steered away from that answer. See Figure 8 in Appendix A.1 for the full results.
@@
These results support the hypothesis that GPT-4 over-relies on the prior probability of potential outputs (see Jang et al. (2023) for some related observations). If the answer has a high probability of occurrence, the model’s priors favor generating it even if its intermediate reasoning steps suggest an alternative output. Conversely, if the answer is of lower likelihood, then even if the chain of reasoning is correct, the priors exert a detrimental influence leading to incorrect final answers.
Prob Chain Steps Output Final Answer
Correct Incorrect
High Correct 19 7
Incorrect 34 40
Low Correct 7 14
Incorrect 1 78
(a) rot-4.
Prob Chain Steps Output Final Answer
Correct Incorrect
High Correct 15 1
Incorrect 55 29
Low Correct 7 9
Incorrect 19 65
(b) rot-13.
Table 1: Confusion matrices (100 examples; 2 probability bins {high, low}) for rot-4 and rot-13. Effect of memorization: incorrect step outputs lead to correct final answers more often for rot-13 than rot-4. Effect of probability: Unfaithfulness has a positive effect more often in high-probability bins than in low-probability bins; having a positive effect means that an incorrect chain is corrected to yield an accurate final answer. Conversely, unfaithfulness has a negative effect more often in low-probability bins than in high-probability bins.
@@ 4.3 Analyzing the effects of noise
The statistically significant impact of $shift\_level$ is evidence that GPT-4’s CoT behavior is in part a noisy version of symbolic reasoning. Accuracy falls as the shift level increases from 1 to 12 and then recovers at higher shift levels (Table 2), consistent with a noisy reasoning process in which deciphering each letter with a shift level of $n$ involves $\min(n,26-n)$ implicit steps, with noise that gives each step some probability of being performed incorrectly. Note that the implicit steps referred to here are different from the steps that are explicitly produced in the chain of thought: the chain of thought uses one explicit step per letter (Figure 2), but here we are discussing the operations that must be implicitly carried out within each step of this chain in order to decode each letter.
@@
We have argued that the relation between accuracy and shift level is evidence that GPT-4 uses a two-way strategy. That is, accuracy is high for small shifts such as 1 but also for large shifts such as 25, which could plausibly be explained by GPT-4 implicitly selecting whichever direction will minimize the number of steps it needs to compute—shifting letters backward for small shift levels or forward for large shift levels. This two-way strategy is effective in that it supports strong performance on large shift levels such as 25. However, we also observe evidence that it contributes to the noise that causes accuracy to decline as the shift level increases. Figure 9 (in Appendix A.4) shows the actual shift level that GPT-4 produces for each letter in its chain of thought, for each of four intended shift levels. Across all four of these cases, GPT-4 shows peaks at both $shift\_level$ and $26-shift\_level$. Thus, while some of the noise affecting the reasoning process may be random, it appears that at least some of the noise can be attributed to confusion between possible shift levels. Decoding a shift level of $n$ can be done by shifting backward $n$ steps or forward $26-n$ steps, but it appears that GPT-4 sometimes mixes up these two strategies by shifting forward $n$ steps (or, equivalently, shifting backward $26-n$ steps), contributing to the overall noise.
@@
Another factor that interacts with noise is temperature. In principle, if CoT entailed pure symbolic reasoning, it would assign 100% probability to the correct continuation (i.e., each predicted next token) and 0% probability to everything else. If so, the temperature used would not affect performance. However, we observe that CoT scores better with a low temperature. For instance, in rot-13, the accuracy is 30.0% at temperature=0 and 0.33% at temperature=1. This shows that its predicted distribution over the vocabulary is not fully discrete—it has some noise in it that a low temperature can remove. However, even a temperature of 0 does not make the performance perfect because noise does not solely arise in the final distribution over the vocabulary (which temperature modifies) but also influences the implicit intermediate steps used to produce that distribution (which temperature does not change).
Style / Shift Level 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25
Text CoT 49.2 52.4 50.8 51.6 42.8 47.0 32.6 44.0 30.2 9.6 0.2 0.0 60.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 16.6 17.0 47.0 45.4 50.4
Math CoT 70.2 68.0 68.6 66.2 58.8 60.4 57.4 63.2 45.4 41.6 40.2 33.8 71.0 16.2 19.2 17.8 11.4 8.8 4.8 1.6 0.8 0.6 0.0 50.0 61.6
Table 2: Accuracy for Text- and Math- CoT prompt styles. Math-CoT performs better than Text-CoT, but both display evidence of memorization because accuracy is highest for shift levels 13, 1, 3, and 2—the most frequent shifts in real-world corpora.
@@ 4.4 Analyzing the effect of memorization
To further investigate memorization, we focus on rot-13, because frequency is generally confounded with simplicity for the other shift levels (e.g., rot-1 is simple as well as frequent). $13$ is the most frequent shift level McCoy et al. (2023), and we observe in Table 2 that GPT-4 shows a spike in accuracy at this shift level in both Text-CoT and Math-CoT, providing strong evidence that memorization plays a role in GPT-4’s CoT performance.
@@ 4.5 The role of intermediate reasoning steps
In both settings, performance was similar to that of standard prompting. This is evidence that CoT depends on “self-conditioning”—explicitly producing text that will be useful as context to condition on when producing the final answer. Merely instructing a model to “think” silently is not helpful, suggesting that the reasoning does not occur internally. These results corroborate prior work finding that CoT is unhelpful when the model is told to produce contentless “filler” tokens instead of contentful text (Lanham et al., 2023); models can reason internally when explicitly trained to do so (Pfau et al., 2024), but current LLMs without this explicit training do not seem to have this ability.
@@ 5 Conclusion
6 Limitations
We have conducted extensive studies on a single task, decoding shift ciphers; we chose this task because it enables us to separate memorization from reasoning in a controlled manner as explained in Subsection 3.1, and these factors cannot be easily disentangled for most other tasks. Our prompts contain one example demonstration (i.e. one-shot CoT prompting), but a single demonstration contains at least 5 reasoning steps which provide more than just one reference of decoding to the model. We also studied a single model, GPT-4, because focusing on just one task and model made it possible to do a deep dive in which we carefully investigated each factor that we considered. An important direction for future work would be to investigate additional tasks and models. In addition, while our experiments showed that GPT-4 displays some hallmarks of true reasoning, they also showed some ways in which it makes errors due to a reliance on memorization and probability; future work could investigate how these limitations could be overcome.
7 Ethical Considerations
@@ Acknowledgements
This work was supported in part by the National Science Foundation SBE Postdoctoral Research Fellowship under Grant No. 2204152. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation
Appendix A Additional Details & Results
A.1 Dataset details
To create our dataset of 7-letter words, we found all 3-letter and 4-letter tokens from the lowercase English alphabet and formed words by considering all possible combinations of a 3-letter token followed by a 4-letter token. Note that GPT-4’s tokenizer distinguishes between word-initial tokens and non-word-initial tokens; thus, to be precise, the tokens that we used were 3-letter word-initial tokens and 4-letter non-word-initial tokens. Following McCoy et al. (2023), we compute the log probability as the log probability that GPT-2 Radford et al. (2019) assigns to the sentence ‘The word is "WORD"’, minus the log probability that it assigns to ‘The word is "’; thus, this yields the log probability assigned to just the word and the following quotation mark in the context of ‘The word is "’. The closing quotation mark is included because it indicates the end of the word. Finally, we manually checked the words in the dataset and filtered them to ensure there were no inappropriate words used. We partitioned each group of 150 examples into two subsets: a subset containing 100 words used to evaluate GPT-4, and a subset containing 50 words used to evaluate logistic regression models that were fitted to GPT-4’s performance on the 100-word subset. We run all evaluations a single time; the accuracies that we report are accuracies over these 100-example sets.
A.2 Standard, Math & Number-CoT prompts
The Standard prompt (Figure 5) used is adopted from McCoy et al. (2023) and yields poor performance on most shift levels. The Number-CoT prompt (Figure 6) in contrast gives nearly perfect scores across the shift levels. It is to be noted that some miscellaneous noise is captured in the Number-CoT case. This arises mostly due to incomplete generations/half-completed chains requiring more tokens than needed as the model does some additional sub-reasoning steps, and in very rare cases produces numbers greater than 25. The Math-CoT prompt used is shown in Figure 7.
Rot-13 text: "fgnl"
Original text: "stay"
shift-13 sequence: "5,6,13,11"
New position = (Given position - 13) mod 26
Using this,
1. 5 -> (5 - 13) mod 26 -> 18
2. 6 -> (6 - 13) mod 26 -> 19
3. 13 -> (13 - 13) mod 26 -> 0
4. 11 -> (11 - 13) mod 26 -> 24
Therefore, the original sequence of numbers is: "18,19,0,24"
shift-13 sequence: <test_input>
Figure 6: The shift cipher task reformatted to be performed in the domain of numbers rather than the domain of letters (Number-CoT). The prompt consists of a description and the demonstration that includes several reasoning steps. Here both the encoded input and decoded output are number sequences abstracting out the alphabet and its influences.
@@ A.2 Standard, Math & Number-CoT prompts
Rot-13 text: <encoded_test_input>
Figure 7: Math-based CoT prompt consisting of a description and the demonstration that includes several reasoning steps.
A.3 Effect of probability
The gap between the overall decoding accuracy and the expected accuracy upon concatenation of step outputs is seen between the orange and blue lines in Figure 8. The drop in accuracy is much higher for the lower probability data (i.e., the blue bin 5 line is lower than the bin 1 line).
Figure 8: Actual overall decoding accuracy vs. faithful accuracy across shift levels showing the flip effects of probability. The effect is amplified for low probability outputs as seen in the larger drop in accuracy between the orange and blue bin 5 (low prob) lines.
A.4 Effect of noise
While decoding can be done by shifting $n$ steps backward or $26-n$ forward, the normalized frequency distribution of the predicted $shift\_level$ in Figure 9 provides evidence that GPT-4 uses a noisy bidirectional reasoning approach by mixing the two strategies—shifting $26-shift\_level$ backward instead of forward.
Figure 9: Normalized frequency distribution vs. predicted $shift\_level$ of step answers for rot-20 to rot-23. The appearance of peaks at $26-shift\_level$ in Math-CoT and Text-CoT prompts is evidence that some of the noise in the model’s predictions is due to confusion between the shift levels $n$ and $26-n$.
A.5 Impact of producing explicit reasoning steps and validity of steps in demonstration.
@@ A.4 Effect of noise
Figure 10: Text-CoT prompt consisting of a description and the demonstration and hidden (left) and partially hidden (right) reasoning steps.
@@
We observe that many of the outputs produced in the case when the reasoning step answers are * are related to the task of shift ciphers. It looks like the model is influenced by terms like ‘cipher’, ‘decode’ that are present in the description; as the reasoning steps no longer help meaningfully as the step output is now *. Influenced by the probabilistic effects discussed before it produces outputs having a high probability and containing words related to security, safety, and programming like ‘encryption’, ‘code cracker’, ‘decoded’, ‘Javascript’, ‘Instagram’.
@@
Figure 11: Prompt snippets with different types of mismatches/corruptions introduced are highlighted in yellow.
With both types of corruption, GPT-4’s performance remained approximately as strong as when it was given accurate, uncorrupted demonstrations. This result corroborates the aforementioned prior work that found no strong correlation between the validity of the reasoning shown in the demonstration and the model performance. The demonstration seems to merely guide the model to solve the task by providing a format to generate accurate reasoning steps Min et al. (2022).

### v2

@@ 4.1 A simple probabilistic approach to modeling the reasoning process
  • $\mathbf{input\_logprob}$: log probability of the encoded input text as measured by GPT-2 Radford et al. (2019). The inputs tend to have a very low probability because they are enciphered.
  • $\mathbf{output\_logprob}$: log probability of the ground-truth output text as measured by GPT-2.
  • $\mathbf{shift\_freq}$: we used the frequency of occurrence of all shift levels that McCoy et al. (2023) provided based on analysis of the C4 corpus Raffel et al. (2020). The assumption is that the distribution of shifts in C4 is similar to the distribution in the training data for GPT-4.
  • $\mathbf{shift\_level}$: the number of steps that must be performed to decode each letter; this feature is added to account for one-way reasoning.
  • $\bm{\min}\mathbf{(shift\_level,26-shift\_level)}$: this value is the minimum number of steps that must be performed to decode each letter, under the assumption that decoding can be done by moving $shift\_level$ steps backward or $(26-shift\_level)$ steps forward; as discussed above, GPT-4 indeed shows evidence of using both of these decoding directions.
@@
$correct\sim input\_logprob+output\_logprob$
$\qquad\qquad+shift\_freq+shift\_level$
$\qquad\qquad+\min(shift\_level,26-shift\_level)$
@@
The following features had a statistically significant effect on model performance: $output\_logprob$, $shift\_freq$, $shift\_level$, and $\min(shift\_level,26-shift\_level)$ ($p<10^{-15}$ in all cases). These results therefore quantitatively support the conclusion that GPT-4 incorporates processes based on probability, memorization, and noisy reasoning (both forward and backward).
@@
Figure 5 shows the predictions of the logistic regression compared to GPT-4’s actual performance. The logistic regression correctly predicts the main trends in GPT-4’s behavior (as expected, it does not match the curve exactly due to the simplicity of the model). In the next few subsections, we conduct some additional experiments that investigate each hypothesized reasoning type in more detail.
Figure 6: Actual overall decoding accuracy vs. faithful accuracy across shift levels showing the effects of probability. The effect is amplified for low probability outputs as seen in the larger drop in accuracy between the orange and blue bin 5 (low probability) lines.
@@ 4.2 Analyzing the effect of probability
If an LLM is influenced by probability, we would expect to occasionally observe unfaithfulness between the chain of reasoning steps produced by the LLM and the LLM’s final answer. Specifically, if the individual reasoning steps would point to a final output that is low-probability, a probabilistic reasoner might instead produce a different final answer that has a higher probability. For example, in our CoT experiments, each step produces one letter, and these letters must be concatenated to form the final answer. If the individual step outputs are S, T, A, Z, the final answer should be STAZ, but a model might instead “self-correct” by producing the higher-probability word STAY.
@@
To investigate unfaithfulness, we compare the faithful accuracy that would be obtained by concatenating GPT-4’s step outputs to the actual overall accuracy. We indeed observe that overall accuracy is generally lower than faithful accuracy, illustrating that unfaithfulness occurs. Further, the drop in accuracy is more pronounced in the low-probability setting than the high-probability setting, which is consistent with the intuition that the lower the probability of a concatenated answer is, the more likely it will be that a probability-reliant model will be steered away from that answer. See Figure 6 for the full results.
@@
These results support the hypothesis that GPT-4 over-relies on the prior probability of potential outputs (see Jang et al. (2023) for some related observations). If the answer has a high probability of occurrence, the model’s priors favor generating it even if its intermediate reasoning steps suggest an alternative output. Conversely, if the answer is of lower probability, then even if the chain of reasoning is correct, the priors exert a detrimental influence leading to incorrect final answers.
Prob Chain Steps Output rot-4 rot-13 Correct Incorrect Correct Incorrect High Correct 19 7 15 1 Incorrect 34 40 55 29 Low Correct 7 14 7 9 Incorrect 1 78 19 65
Table 1: Confusion matrices (100 examples; 2 probability bins {high, low}) for rot-4 and rot-13. Effect of memorization: incorrect step outputs lead to correct final answers more often for rot-13 than rot-4. Effect of probability: Unfaithfulness has a positive effect more often in high-probability bins than in low-probability bins; Conversely, unfaithfulness has a negative effect more often in low-probability bins than in high-probability bins.
@@ 4.3 Analyzing the effects of noise
The statistically significant impact of $shift\_level$ is evidence that GPT-4’s CoT behavior is in part a noisy version of symbolic reasoning. Accuracy falls as the shift level increases from 1 to 12 and then recovers at higher shift levels (Figure 8), consistent with a noisy reasoning process in which deciphering each letter with a shift level of $n$ involves $\min(n,26-n)$ implicit steps, with noise that gives each step some probability of being performed incorrectly. Note that the implicit steps referred to here are different from the steps that are explicitly produced in the chain of thought: the chain of thought uses one explicit step per letter (Figure 3), but here we are discussing the operations that must be implicitly carried out within each step of this chain in order to decode each letter.
Figure 7: Normalized frequency distribution vs. predicted $shift\_level$ of step answers for rot-20 to rot-23. The appearance of peaks at $26-shift\_level$ in Math-CoT and Text-CoT prompts showcases the model’s noisy attempt in taking the shorter path—i.e., moving $26-x$ shifts forward.
@@
We have argued that the relation between accuracy and shift level is evidence that GPT-4 uses a two-way strategy. That is, accuracy is high for small shifts such as 1 but also for large shifts such as 25, which could plausibly be explained by GPT-4 implicitly selecting whichever direction will minimize the number of steps it needs to compute—shifting letters backward for small shift levels or forward for large shift levels. This two-way strategy is effective in that it supports strong performance on large shift levels such as 25. However, we also observe evidence that it contributes to the noise that causes accuracy to decline as the shift level increases. Figure 7 shows the actual shift level that GPT-4 produces for each letter in its chain of thought, for each of four intended shift levels. Across all four of these cases, GPT-4 shows peaks at both $shift\_level$ and $26-shift\_level$. Thus, while some of the noise affecting the reasoning process may be random, it appears that at least some of the noise can be attributed to confusion between possible shift levels. Decoding a shift level of $n$ can be done by shifting backward $n$ steps or forward $26-n$ steps, but it appears that GPT-4 sometimes mixes up these two strategies by shifting forward $n$ steps (or, equivalently, shifting backward $26-n$ steps), contributing to the overall noise.
@@
Another factor that interacts with noise is temperature. In principle, if CoT entailed pure symbolic reasoning, it would assign 100% probability to the correct continuation (i.e., each predicted next token) and 0% probability to everything else. If so, the temperature used would not affect performance. However, we observe that CoT scores better with a low temperature. For instance, in rot-13, GPT-4’s accuracy is 30.0% at temperature=0 and 0.33% at temperature=1. This shows that its predicted distribution over the vocabulary is not fully discrete—it has some noise in it that a low temperature can remove. However, even a temperature of 0 does not make the performance perfect because noise does not solely arise in the final distribution over the vocabulary (which temperature modifies) but also influences the implicit intermediate steps used to produce that distribution (which temperature does not change).
Figure 8: Accuracy for Text- and Math- CoT prompt styles with GPT-4. Math-CoT performs better than Text-CoT, but both display evidence of memorization as accuracy is highest for shift level 13—the most frequent shift in real-world corpora.
@@ 4.4 Analyzing the effect of memorization
To further investigate memorization, we focus on rot-13, because frequency is generally confounded with simplicity for the other shift levels (e.g., rot-1 is simple as well as frequent). $13$ is the most frequent shift level McCoy et al. (2023), and we observe in Figure 8 that GPT-4 shows a spike in accuracy at this shift level in both Text-CoT and Math-CoT, providing strong evidence that memorization plays a role in GPT-4’s CoT performance.
@@ 4.5 The role of intermediate reasoning steps
In both settings, performance was similar to that of the standard prompting variant (shown in Figure 1, panel 2). This is evidence that CoT depends on “self-conditioning”—explicitly producing text that will be useful as context to condition on when producing the final answer. Merely instructing a model to “think” silently is not helpful, suggesting that the reasoning does not occur internally. These results corroborate prior work finding that CoT is unhelpful when the model is told to produce contentless “filler” tokens instead of contentful text (Lanham et al., 2023); models can reason internally when explicitly trained to do so (Pfau et al., 2024), but current LLMs without this explicit training do not seem to have this ability.
@@ 5 Conclusion
Limitations
We have conducted extensive studies on a single task, decoding shift ciphers; we chose this task because it enables us to separate memorization from reasoning in a controlled manner as explained in Subsection 3.1, and these factors cannot be easily disentangled for most other tasks. Our prompts contain one example demonstration (i.e. one-shot CoT prompting), but a single demonstration contains multiple reasoning steps which provide more than just one reference of decoding to the model. In addition, while our experiments showed that these frontier models display some hallmarks of true reasoning, they also showed some ways in which they make errors due to a reliance on memorization and probability; future work could investigate how these limitations could be overcome.
Ethical Considerations
@@ Acknowledgements
This work was supported in part by the National Science Foundation SBE Postdoctoral Research Fellowship under Grant No. 2204152. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
Appendix A Additional Details & Experiments
A.1 Math- & Number- CoT prompts
While the Standard prompt (Figure 2) yields poor performance on most shift levels, the Number-CoT prompt (Figure 10) in contrast gives nearly perfect scores across the shift levels. It is to be noted that some miscellaneous noise is captured in the Number-CoT case. This arises mostly due to incomplete generations/half-completed chains requiring more tokens than needed as the model does some additional sub-reasoning steps, and in very rare cases produces numbers greater than 25. The Math-CoT prompt used is shown in Figure 9.
@@ A.1 Math- & Number- CoT prompts
Rot-13 text: <test_input>
Figure 9: Math-CoT prompt.
shift-13 sequence: "5,6,13,11"
New position = (Given position - 13) mod 26
Using this,
1. 5 -> (5 - 13) mod 26 -> 18
2. 6 -> (6 - 13) mod 26 -> 19
3. 13 -> (13 - 13) mod 26 -> 0
4. 11 -> (11 - 13) mod 26 -> 24
Therefore, the original sequence of numbers is: "18,19,0,24"
shift-13 sequence: <test_input>
Figure 10: The shift cipher task reformatted to be performed in the domain of numbers rather than the domain of letters (Number-CoT). Here both the encoded input and decoded output are number sequences abstracting out the alphabet and its influences.
Figure 11: Accuracies with Claude 3 and Llama 3.1. The top panel shows the performance when prompted with Standard prompting. The bottom panel shows the trend with Text-CoT. The shaded regions indicate the gap between the low and high probability bins.
A.2 Results with Llama 3.1 and Claude 3
We display the results with Llama 3.1 and Claude 3 when prompted with standard prompts and Text-CoT in Figure 11. Interestingly, Llama 3.1 was trained on CoT data, such that even when it is prompted with standard prompts it generates reasoning steps. However, as we can see from Figure 11, with Text-CoT there is a further enhancement in scores across the shift levels. Overall, both models display trends similar to that of GPT-4 shown in Figure 1.
A.3 Impact of producing explicit reasoning steps and validity of steps in demonstration.
@@
Figure 12: Text-CoT prompt consisting of hidden (top) and partially hidden (bottom) reasoning steps.
@@
We observe that many of the outputs produced in the case when the reasoning step answers are * are related to the task of shift ciphers. It seems that the model now relies on terms such as cipher and decode that are present in the description, perhaps relying less on the reasoning steps because these reasoning steps are now less informative given that each step’s output is now *. Influenced by the probabilistic effects discussed before it produces outputs having a high probability and containing words related to security, safety, and programming such as encryption, code cracker, decoded, Javascript, and Instagram.
@@
Figure 13: Prompt snippets with different types of mismatches/corruptions. The corruptions that were introduced are highlighted in yellow.
With both corruptions, GPT-4’s performance remained approximately as strong as when it was given accurate, uncorrupted demonstrations. This result corroborates the aforementioned prior work that found no strong correlation between the validity of the reasoning shown in the demonstration and the model performance. The demonstration seems to merely guide the model to solve the task by providing a format to generate accurate reasoning steps Min et al. (2022).